OpenStudy (anonymous):

What is the end behavior of the function f(x) = 3x4-x3 + 2x2 + 4x + 5?

OpenStudy (whpalmer4):

@larryboxaplenty no, it's not a parabola...a parabola would have the form \(f(x) = ax^2+bx+c\) end behavior depends only on the highest order term, here \(3x^4\). What happens to the value of \(3x^4\) as \(x\rightarrow \infty\)? What about as \(x\rightarrow -\infty\)?